# ND-SDF: Learning Normal Deflection Fields for High-Fidelity Indoor Reconstruction

### v1

@@ 4 Experiment
Ours 0.032 0.044 84.9 73.5 0.038 78.6
Table 1: Quantitative results on the ScanNet dataset. We achieve state-of-the-art performance.
Datasets We conducted experiments on four indoor datasets: ScanNet Dai et al. (2017), Replica Straub et al. (2019), TanksandTemples Knapitsch et al. (2017), and ScanNet++ Yeshwanth et al. (2023). ScanNet contains 1513 indoor scenes captured with an iPad, processed using the BundleFusion algorithm to obtain camera poses and surface reconstruction. Replica is a synthetic dataset comprising 18 indoor scenes. Each scene features dense geometry and high dynamic range textures. TanksandTemples is a large-scale 3D reconstruction dataset including high-resolution outdoor and indoor environments. We followed the splits from MonoSDF and applied the same evaluation settings. We also conduct experiments on ScanNet++, which includes 460 indoor scenes captured using laser scanners These scenes offer high-quality dense reconstructions and images. For testing, we selected six scenes from ScanNet++.
Implementation details Our method was implemented using PyTorch Paszke et al. (2019). The image resolution for all scenes is 384×384. We obtained normal and depth cues using Omnidata. Multi-resolution hash grids were utilized for scene representation. Both the geometry network and color network consisted of two layers, each with 256 nodes. Our network was optimized using AdamW Loshchilov & Hutter (2017) optimizer with a learning rate of 1e-3. The weights for loss terms were: $\lambda_{1}=0.05$, $\lambda_{2}=0.0005$, $\lambda_{3}=0.025$, $\lambda_{4}=0.05$. Upon adequate initialization of the deflection field, we initiated deflection angle guided sampling, photometric optimization, and unbiased rendering. All experiments were conducted on an NVIDIA TESLA A100 PCIe 40GB, with each iteration sampling 4$\times$1024 rays, totaling 128,000 training steps. The total training time was about 12 hours. Our hash encoding resolution spanned from $2^{5}$ to $2^{11}$ across 16 levels, with the initial activation level set to 8 and activation steps to 2000.
Metrics In line with prior research, we employ six standard metrics to evaluate the reconstructed meshes: Accuracy, Completeness, Chamfer Distance, Precision, Recall, and F-score. Additionally, normal consistency is utilized for evaluating the Replica dataset.
Baselines We compare with the following methods: (1) Classic MVS COLMAP Schonberger & Frahm (2016); Schönberger et al. (2016); (2) Neural implicit methods including VolSDF and Neus; (3) Methods utilizing auxiliary data including MonoSDF and NeuRIS; (4) Other indoor reconstruction methods like HelixSurf Liang et al. (2023).
Figure 3: Qualitive results on ScanNet. Compared with previous state-of-the-art works, our method produces more thin and fine-grained structures, such as the chair legs and detailed table layouts.
@@
Table 2: F-score on the TanksandTemples dataset.
@@ 4.1 Comparisons
ScanNet: In Figure 3, we present qualitative results and in Table 1, quantitative results are displayed. These quantitative results are averaged over the selected four scenes. ND-SDF surpasses previous best-performing works, achieving state-of-the-art performance. Specifically, we achieve the highest F-score, which serves as a credible indicator reflecting reconstruction accuracy by considering both Accuracy and Completeness. When compared with HelixSurf and MonoSDF (as shown in Figure 3), our approach accurately captures thin structures such as chair legs and lamp rings. This underscores the superiority of our proposed deflection methods, which significantly enhance the recovery of thin and fine-grained structures indoors.
@@
Table 3: Quantitative results on the ScanNet++ dataset. We select 6 scenes and calculate the average metrics of them.
@@
Table 4: Quantitative results of Replica dataset.
Replica: Following MonoSDF Yu et al. (2022b), we used their pre-processed masks to filter out abnormal priors. We reported quantitative results in Table 4. Our approach achieves the highest F-score and the lowest Chamfer distance, substantially surpassing previous works. See supplementary for qualitative results.
Tanks and temples: We also conducted experiments on the advanced split of T&T, which provides several challenging large-scale indoor scenes. Quantitative results are provided in Table 2, where ND-SDF outperforms previous methods that utilize priors. Additional qualitative results can be found in the supplementary material, our method recovered substantial detailed structures.
@@
Table 5: F-score of Different Priors Models..
ScanNet++: ScanNet++ comprises numerous complex indoor scenes, each offering high-resolution GT mesh and RGB data. Since no public results are available for ScanNet++, we trained some baselines using our settings, including VolSDF Yariv et al. (2021), MonoSDF Yu et al. (2022b), and Baked-Angelo Yu et al. (2022a). Among these, only MonoSDF utilized monocular priors. Our quantitative results are summarized in Table 3, demonstrating state-of-the-art performance across all metrics. For additional visualization results, please refer to the supplementary material.
@@
Table 6: Quantitative Results of Ablation for Different Modules. ‘Base’ can be considered as MonoSDF+Neuralangelo. ‘ND’ denotes applying the deflection field and the deflected normal loss term $L^{d}_{normal}$. ‘AP’ denotes the two adaptive deflection angle prior loss terms including $L^{ad}_{normal}$ and $L^{ad}_{depth}$. ‘DO’ denotes the two deflection angle guided optimizations, namely sampling and photometric loss. ‘DU’ denotes the deflection angle guided unbiased rendering.
Figure 4: Visualization Results of the Models after Ablation.
@@ 4.2.1 Ablation of different modules
Analysis of Normal Deflection Field The Base model strictly adheres to monocular cues, leading to significant detail loss due to misguided priors in complex areas. In ModelA, we introduced the deflection field and applied the deflected normal loss to learn deviations. The improvement in the F-score to 0.632 (as shown in Table 6) confirms the superiority of the deviation learning approach. However, we observed wrinkles (as depicted in Figure 4) in textureless regions like walls. This issue arises from the lack of constraints on deflection, resulting in inefficient utilization of priors in smooth areas. In ModelB, introducing an adaptive deflection angle prior loss significantly enhanced reconstruction quality, balancing both smoothness and detail.
Analysis of deflection angle guided optimization While ModelB substantially improves reconstruction quality compared to the Base, it still struggles with finer structures. In ModelC, we introduced deflection angle guided sampling and color loss to emphasize surface details. Finally, we incorporated deflection angle guided unbiased rendering to facilitate the recovery of thin structures. Figure 4 demonstrates that Ours recovers a wide range of complex and fine structures, including skull models, wires, and iron swabs, surpassing the Base model. As indicated in Table 6, Ours achieved the highest f-score, proving the effectiveness of the proposed modules for restoring high-fidelity surfaces.
@@ 4.2.2 Ablation of different prior models
Our method learns deviations between the scene and normal priors predicted by a specific pretrained model. It is essential to explore its effectiveness across different cues. To achieve this, we employed the currently state-of-the-art method Geowizard Fu et al. (2024) to generate both depth and normal cues. The quantitative results, as summarized in Table 5, demonstrate that our method significantly enhances surface quality regardless of the cues used. This experiment demonstrates the superiority of our method. It invariably improves reconstruction accuracy in scenarios where prior knowledge is used.
@@ 5 Conclusion
We have presented ND-SDF, a novel approach that learns deviations between the scene and normal priors for high-fidelity indoor surface reconstruction. We propose an adaptive deflection angle prior loss to dynamically supervise areas with varying characteristics. By identifying high-frequency regions based on deflection angles, we employ angle guided optimization to generate thin and fine-grained structures. Our method recovers a substantial amount of complex structures, as demonstrated by extensive qualitative and quantitative results that confirm its superiority. ND-SDF efficiently adapts to unreliable priors in challenging areas, significantly addressing the issue of detail loss caused by large prior errors for regions with fine structures.
Limitation ND-SDF is confined to reconstruction scenarios involving priors. Our approach exhibits limited capability in handling areas with less coverage. Due to inherent ambiguities in observations, the deflection field may learn incorrect structures and settle into local optima. We kindly refer the reader to the supplementary material for more analysis and results. In the future, we may explore multi-view consistency constraints to enhance reconstruction quality.
Appendix / supplemental material
Figure 5: Architecture of ND-SDF. We utlize the multi-resolution hash grids $\gamma_{L}$ as the scene representation. The geometry network $f_{g}$, color network $f_{c}$, and deflection network $f_{d}$ are all constructed using simple multilayer perceptrons (MLPs).
@@ A.1 Architecture
When considering the inputs to the deflection network, two scenarios are contemplated: (1) No deflection is required when the prior is entirely accurate. In this case, the quaternion axis aligns with the scene normals, thereby closely associating deflection with the scene geometry. (2) Given that the prior model is view-dependent, the quaternion for alignment with the prior must also be view-dependent. Consequently, we define:
@@
We observe that the structure of the deflection network mirrors that of the color network Yariv et al. (2020), denoted as $\textbf{c}=f_{c}(\textbf{x},\textbf{v},\textbf{n},\textbf{z})$. However, considering that the color network solely models scene-related radiance, while the deflection network encapsulates a greater extent of the prior deviation, we separate these two networks.
@@ A.2 Progressive warm-up
In the early stages of training, deflection fields would introduce negative effects such as noise. Before the rough formation of scene geometry, these deflection quaternions are random and erroneous. If unconstrained, deflecting the true normals during this initial phase can lead to reconstruction errors, such as surface protrusions.
@@
$\begin{array}[]{l}\frac{\theta}{2}=\arccos{\textbf{Q}_{0}}\\ \textbf{u}=\frac{[\textbf{Q}_{1},\textbf{Q}_{2},\textbf{Q}_{3}]}{\sin{\frac{\theta}{2}}}\end{array}.$ (18)
@@
$\begin{array}[]{cc}\frac{\theta_{iter}}{2}=prog_{q}\times\frac{\theta}{2}\\ \textbf{u}_{iter}=prog_{q}\times\textbf{u}+(1-prog_{q})\times\hat{\textbf{N}}(\textbf{r})\end{array}.$ (19)
@@
During the initial training stages, the deflection angles are close to zero, enabling us to leverage prior knowledge to learn smooth regions and approximate the scene geometry. We consider it to be adequately initialized as the warm-up phase concludes. In this process, the deflection field gradually adapts to more complex structures and deviations from the normal priors. The quantitative results with and without progressive warm-up are presented in Table 7.
Method F-score
w/o warm-up 0.667
w warm-up 0.686
Table 7: Ablation of Warm-Up Strategy. The warm-up strategy can better facilitate the initialization of the deflection field, resulting in improved surface reconstruction quality.
A.3 More details
@@ A.3.1 Adaptive deflection angle prior loss
In Section 3.3, we introduce two modulation functions, namely $g^{d}$ and $g$ (Eq. 12), which adjust the weight of the deflected and original normal loss terms based on the deflection angle.
Here provide the specific form for both functions. For convenience, we employ a shifted logistic function to define them:
$\begin{array}[]{cc}g^{d}(\theta)=\frac{1}{(1+e^{-s_{0}}(\theta-\theta_{0}))}\\ g(\theta)=1-g^{d}(\theta)\end{array}.$ (21)
@@
As illustrated in Figure 6, we consistently observe that a deviation angle less than 5° indicates a simple flat region, whereas a deviation angle greater than 15° largely suggests a complex structural region.
Additionally, we visualize the modulation function $g^{d}(\Delta\theta)$ in Figure 7. For regions with a deviation angle less than 5°, we assign weights greater than 0.9 to the original normal loss term to effectively utilize the normal priors in smooth regions.
Figure 7: Modulation Function $g^{d}(\Delta\theta)$ defined in Eq. 21.
@@ A.3.2 Deflection angle guided optimization
To achieve ray sampling guided by the deflection angle, we dynamically maintain a deflection angle map ($\Delta\bar{\theta}$) per image during the training process. In the sampling phase, we first calculate the per-pixel sampling probability based on the deflection angle map. Then use the probability map for inverse sampling, meaning that areas with greater deflection magnitude tend to have more rays sampled.
@@
$p(\textbf{r}_{i})=1+\frac{t_{1}}{(1+e^{-s_{1}}(\Delta\bar{\theta}(\textbf{r}_{i})-\theta_{1}))},$ (22)
where $\Delta\bar{\theta}$ represents the deflection angle map, and $p(\textbf{r}_{i})\in[1,t_{1}+1]$, with the steepness controlled by $s_{1}$. We then normalize the probability map for inverse sampling using the following equation:
@@
For the angle guided color loss, we similarly employ a scaled and shifted logistic function to re-weight the original color loss term guided by the deflection angle (Eq. 14):
$w_{color}(\Delta\theta(\textbf{r}))=1+\frac{t_{2}}{(1+e^{-s_{2}}(\Delta{\theta}(\textbf{r})-\theta_{2}))}.$ (24)
Regarding the angle guided unbiased rendering, we define a bias confidence using the deflection angle. Another shifted logistic function is employed to compute the bias confidence:
$cfd({\textbf{r}})=\frac{1}{(1+e^{-s_{3}}(\Delta\bar{\theta}(\textbf{r})-\theta_{3}))}.$ (25)
This leads to deflection angle guided unbiasing:
$\sigma(\textbf{r}({t_{i}}))=\alpha\Psi_{\beta}\left(\frac{-f_{g}(\textbf{r}({t_{i}}))}{cfd(\textbf{r})\left|f_{g}^{\prime}(\textbf{r}({t_{i}}))\right|+1-cfd(\textbf{r})}\right).$ (26)
In practice, we set: $(s_{1}=25,\theta_{1}=\frac{\pi}{12},t_{1}=4),(s_{2}=25,\theta_{2}=\frac{\pi}{12},t_{2}=2),(s_{3}=25,\theta_{3}=\frac{\pi}{18})$.

### v2

@@ 4 Experiment
DebSDF Xiao et al. (2024) 0.036 0.040 80.7 76.5 0.038 78.5
H2OSDF Park et al. (2024) 0.032 0.037 83.4 76.9 0.035 79.9
Ours 0.031 0.036 84.0 80.3 0.034 82.0
Table 1: Quantitative results on the ScanNet dataset. Our method substantially outperforms previously best-performing methods across all metrics to date.
Datasets We conducted experiments on four indoor datasets: ScanNet Dai et al. (2017), Replica Straub et al. (2019), TanksandTemples Knapitsch et al. (2017), and ScanNet++ Yeshwanth et al. (2023). ScanNet contains 1513 indoor scenes captured with an iPad, processed using the BundleFusion algorithm to obtain camera poses and surface reconstruction. Replica is a synthetic dataset comprising 18 indoor scenes. Each scene features dense geometry and high dynamic range textures. TanksandTemples is a large-scale 3D reconstruction dataset including high-resolution outdoor and indoor environments. We followed the splits from MonoSDF and applied the same evaluation settings. We also conducted experiments on ScanNet++, which includes 460 indoor scenes captured using laser scanners. These scenes offer high-quality dense reconstructions and images. For testing, we selected six scenes from ScanNet++.
Implementation details Our method was implemented using PyTorch Paszke et al. (2019). The image resolution for all scenes is 384×384. We obtained normal and depth cues using Omnidata Eftekhar et al. (2021). Multi-resolution hash grids were utilized for scene representation. Both the geometry network and color network consists of two layers, each with 256 nodes. Our network was optimized using AdamW Loshchilov & Hutter (2017) optimizer with a learning rate of 1e-3. The weights for loss terms were: $\lambda_{1}=0.05$, $\lambda_{2}=0.0005$, $\lambda_{3}=0.025$, $\lambda_{4}=0.05$. Upon adequate initialization of the deflection field, we initiated deflection angle guided sampling, photometric optimization, and unbiased rendering. All experiments were conducted on an NVIDIA TESLA A100 PCIe 40GB, with each iteration sampling 4$\times$1024 rays, totaling 128,000 training steps. The training time was about 14 GPU hours. Our hash encoding resolution spanned from $2^{5}$ to $2^{11}$ across 16 levels, with the initial activation level set to 8 and activation steps set to 2000.
Metrics In accordance with prior researches, we employ six standard metrics to evaluate the reconstructed meshes: Accuracy, Completeness, Chamfer Distance, Precision, Recall, and F-score. Additionally, normal consistency is used to evaluate the Replica dataset.
Baselines We compare our approach with the following methods: (1) traditional MVS techniques COLMAP Schonberger & Frahm (2016); Schönberger et al. (2016); (2) basic neural implicit methods, including VolSDF and NeuS; (3) methods incorporating auxiliary data, including MonoSDF and NeuRIS; (4) other indoor reconstruction methods, such as HelixSurf Liang et al. (2023).
Figure 4: Qualitative results on ScanNet. Compared to previous state-of-the-art methods, our approach generates more thin and fine-grained structures, such as chair legs and detailed table layouts.
@@
Table 2: F-score on the TanksandTemples dataset, evaluated at a resolution of 768.
@@ 4.1 Comparisons
ScanNet: In Figure 4, we present qualitative results and in Table 1, quantitative results are displayed. These quantitative results are averaged over the selected four scenes. ND-SDF significantly surpasses previous best-performing works, achieving state-of-the-art performance. Specifically, we achieve the highest F-score, a reliable metric that reflects reconstruction accuracy by balancing both Accuracy and Completeness. When compared with HelixSurf and MonoSDF (as shown in Figure 4), our approach accurately captures thin structures, such as chair legs and lamp rings. The results highlight the effectiveness of our proposed deflection methods, which substantially enhance the surface accuracy and the recovery of thin and fine-grained structures in indoor environments.
@@
Table 3: Quantitative results on the ScanNet++ dataset. We select 6 scenes and calculate their average metrics. Baked-Angelo is the best-performing reconstruction method tested without priors.
@@
Table 4: Quantitative results of Replica dataset, significantly surpassing previous methods.
Replica: Following MonoSDF Yu et al. (2022b), we used their pre-processed masks to filter out abnormal priors. Quantitative results are presented in Table 4, where our approach achieves the highest F-score and the lowest Chamfer Distance, significantly surpassing previous methods. Refer to the supplementary material for visualization results.
Tanks and temples: We also conducted experiments on the advanced split of T&T, which features several challenging large-scale indoor scenes. As shown in Table 2, ND-SDF surpasses previous methods that utilize priors in terms of quantitative performance. Additional qualitative results can be found in the supplementary material, where our method recovered substantial detailed structures.
@@
Table 5: F-score of Different Priors Models.
ScanNet++: ScanNet++ comprises numerous complex indoor scenes, each offering high-resolution GT mesh and RGB data. Since no public results are available for ScanNet++, we trained several baselines methods, including VolSDF, MonoSDF, and Baked-Angelo Yu et al. (2022a). Among these, only MonoSDF utilized monocular priors as we did. The quantitative results are summarized in Table 3, demonstrating the state-of-the-art performance of our method across all metrics. For additional visualization results, please refer to the appendix.
@@
Table 6: Quantitative Results of Ablation for Different Modules. ‘Base’ can be viewed as a simple combination of MonoSDF and Neuralangelo. ‘ND’ denotes applying the deflection field and the deflected normal loss term $L^{d}_{normal}$. ‘AP’ refers to the adaptive deflection angle prior loss terms including $L^{ad}_{normal}$ and $L^{ad}_{depth}$. ‘DO’ denotes the two deflection angle guided optimizations, namely sampling and photometric loss. ‘DU’ denotes the deflection angle guided unbiased rendering.
Figure 5: Visualization Results of the Models after Ablation. Detailed definitions of these models are provided in Table 6.
@@ 4.2.1 Ablation of different modules
Analysis of Normal Deflection Field The Base model strictly adheres to monocular cues, leading to significant detail loss due to misguided priors in complex areas. In ModelA, we introduced the deflection field and directly applied the deflected normal loss to learn deviations. The improvement in the F-score to 0.632 (as shown in Table 6) confirms the superiority of the deviation learning approach. However, we observed wrinkles (as depicted in Figure 5) in textureless regions like walls. This issue arises from the lack of constraints on the deflection, resulting in inefficient utilization of priors in smooth regions. In Model B, the introduction of the adaptive deflection angle prior loss significantly enhances reconstruction quality, striking a balance between smoothness and detail.
Analysis of deflection angle guided optimization While ModelB substantially improves reconstruction quality compared to the Base, it still struggles with finer structures (as depicted in Figure 5). In ModelC, we introduced deflection angle guided sampling and color loss to emphasize surface details. Finally, we incorporated an adaptive unbiased rendering method to facilitate the recovery of thin structures. Figure 5 illustrates that Ours recovers a wide range of complex and fine structures, such as skull models and iron swabs, surpassing the Base model. As indicated in Table 6, Ours achieved the highest F-score, quantitatively outperforming other models. This ablation study strongly validates the effectiveness of the proposed modules in restoring high-fidelity surfaces.
@@ 4.2.2 Ablation of different prior models
Our method learns deviations between the scene and normal priors predicted by a specific pretrained model. It is essential to evaluate its effectiveness across cues generated by different pretrained models. To achieve this, we employed the state-of-the-art method Geowizard Fu et al. (2024) to generate both depth and normal cues. The quantitative results, summarized in Table 5, demonstrate that our method significantly enhances surface quality irrespective of the cues utilized. This experiment illustrates the superiority of our approach. It invariably improves reconstruction accuracy in scenarios where prior knowledge is used.
@@ 5 Conclusion
We have presented ND-SDF, a novel approach that learns deviations between the scene and normal priors for high-fidelity indoor surface reconstruction. We introduce an adaptive deflection angle prior loss to dynamically supervise areas with varying characteristics. By identifying high-frequency regions based on deflection angles, we employ angle guided optimization to generate thin and fine-grained structures. Our method recovers a substantial number of complex structures, as demonstrated by extensive qualitative and quantitative results that confirm its superiority. ND-SDF efficiently adapts to unreliable priors in challenging areas, significantly addressing the detail loss issue caused by large prior errors in regions with fine structures.
Limitation ND-SDF is confined to reconstruction scenarios involving priors. Our approach exhibits limited capability in handling areas with less coverage. Due to inherent ambiguities in observations, the deflection field may learn incorrect structures and converge to local optima. We encourage the reader to refer the supplementary material for more analysis and results. In the future, we may explore multi-view consistency constraints to enhance reconstruction quality.
Appendix / supplementary material
Figure 6: Architecture of ND-SDF. We utlize multi-resolution hash grids $\gamma_{L}$ for scene representation. The geometry network $f_{g}$, color network $f_{c}$, and deflection network $f_{d}$ are all constructed using simple multilayer perceptrons (MLPs).
@@ A.1 Architecture
When considering the inputs to the deflection network, two scenarios are contemplated: (1) No deflection is required when the prior is completely accurate. In this case, the quaternion axis aligns with the scene normals, indicating that the deflection axis is closely related to the scene geometry; (2) Since the prior model is view-dependent, the quaternion for alignment with the prior must also be view-dependent. Consequently, we define:
@@
We observe that the structure of the deflection network mirrors that of the color network Yariv et al. (2020), denoted as $\textbf{c}=f_{c}(\textbf{x},\textbf{v},\textbf{n},\textbf{z})$. However, considering that the color network exclusively models scene-related radiance, while the deflection network encapsulates a broader range of prior deviations, we separate these two networks.
@@ A.2 Progressive warm-up
In the early stages of training, deflection fields may introduce negative effects, such as noise. Before the rough formation of scene geometry, these deflection quaternions are random and erroneous. If unconstrained, deflecting the true normals during this initial phase can lead to reconstruction errors, such as surface protrusions.
@@
$\frac{\theta}{2}&=\arccos{\textbf{Q}_{0}}\\ \textbf{u}&=\frac{[\textbf{Q}_{1},\textbf{Q}_{2},\textbf{Q}_{3}]}{\sin{\frac{\theta}{2}}}\end{aligned}.\right.$ (18)
@@
$\frac{\theta_{iter}}{2}&=prog_{q}\times\frac{\theta}{2}\\ \textbf{u}_{iter}&=prog_{q}\times\textbf{u}+(1-prog_{q})\times\hat{\textbf{N}}(\textbf{r})\end{aligned}.\right.$ (19)
@@
During the initial training stage, the deflection angles start at zero and gradually increase, allowing the model to effectively leverage prior knowledge to optimize smooth regions and approximate the scene geometry. We consider it to be adequately initialized upon the warm-up phase concludes. Throughout this process, the deflection field gradually adapts to more complex structures and learns the deviations between the true geometry and the normal priors. Quantitative results with and without progressive warm-up are presented in Table 7.
Method w/o warm-up w warm-up
F-score 0.667 0.686
Table 7: Ablation of Warm-Up Strategy on ScanNet++. The warm-up strategy can better facilitate the initialization of the deflection field, resulting in improved surface reconstruction quality.
A.3 Detailed Implementation Aspects
@@ A.3.1 Adaptive deflection angle prior loss
In Section 3.3, we introduce two modulation functions, namely $g^{d}$ and $g$ (Eq. 12), which adjust the weights of the deflected and original normal loss terms based on the deflection angle.
Here provide the specific forms for both functions. We employ a shifted logistic function to define them:
[…]
@@
As shown in Figure 7, a deviation angle of less than 5° consistently indicates a simple flat region, whereas an angle greater than 15° typically denotes a complex structural region.
Additionally, we visualize the modulation function $g^{d}(\Delta\theta)$ in Figure 8, with $(s_{0},\theta_{0})$ set to $(12.5,\frac{\pi}{12})$. For regions with a deviation angle less than 5°, weights greater than 0.9 are assigned to the original normal loss term, ensuring the effective incorporation of normal priors in smooth regions.
To further demonstrate the effectiveness of the proposed adaptive deflection angle prior loss, we visualize three normal loss heatmaps: the naive normal loss, the deflected normal loss, and the adaptive deflection angle normal loss (as shown in Figure 9). The black box highlights thin structures and ambiguous regions where monocular priors are largely unreliable. In these areas, the naive normal loss heatmap shows significant errors, while both the deflected and adaptive losses exhibit significantly lower errors by aligning the deflected scene normals with the normal priors (GT). In summary, the adaptive prior loss can dynamically adjust the weighting between the naive and deflected loss terms, promoting globally correct convergence while fully leveraging the normal prior.
Figure 8: Modulation Functions defined in Eq. 21.
Figure 9: Heatmap comparison of naive, deflected, and adaptive deflection angle normal loss.
@@ A.3.2 Deflection angle guided optimization
To implement ray sampling guided by the deflection angle, we dynamically maintain a deflection angle map ($\Delta\bar{\theta}$) for each image during the training process. We update the angle map using the following equation:
$\Delta\bar{\theta}(\textbf{r})=\max{(\Delta\bar{\theta}^{old}(\textbf{r})\times\eta,\Delta\theta(r))},$ (22)
where r is the sampled ray and $\eta$ is the decay coefficient. We initialize the angle map to zero at the beginning.
In the sampling phase, we first calculate the per-pixel sampling probability based on the angle map, then use this probability map for inverse sampling, such that areas with greater deflection magnitude tend to have more rays sampled.
@@
$p(\textbf{r}_{i})=1+\frac{t_{1}}{1+e^{-s_{1}}(\Delta\bar{\theta}(\textbf{r}_{i})-\theta_{1})},$ (23)
where $\Delta\bar{\theta}$ represents the maintained angle map, and $p(\textbf{r}_{i})\in[1,t_{1}+1]$, with the steepness controlled by $s_{1}$. We then normalize the probability map for inverse sampling using the following equation:
@@
For the angle guided color loss, we similarly employ a scaled and shifted logistic function to re-weight the original color loss term based on the deflection angle (Eq. 14):
$w_{color}(\Delta\theta(\textbf{r}))=1+\frac{t_{2}}{1+e^{-s_{2}}(\Delta{\theta}(\textbf{r})-\theta_{2})}.$ (25)
Regarding the unbiased rendering guided by the deflection angle, we first define a bias confidence using another variant of the logistic function to compute it:
$cfd({\textbf{r}})=\frac{1}{1+e^{-s_{3}}(\Delta\bar{\theta}(\textbf{r})-\theta_{3})}.$ (26)
Note that we utilize the maintained angle map to calculate the confidence instead of the deflection angle $\Delta\theta(\textbf{r})$ computed from current iteration, as unbiased densities of the sampled points along the ray are required prior to the volume compositing process. This then leads to the deflection angle guided unbiasing:
$\sigma(\textbf{r}({t_{i}}))=\frac{1}{\beta}\Psi_{\beta}\left(\frac{-f_{g}(\textbf{r}({t_{i}}))}{cfd(\textbf{r})\left|f_{g}^{\prime}(\textbf{r}({t_{i}}))\right|+1-cfd(\textbf{r})}\right).$ (27)
In practice, we set the parameters as follows: $(s_{1}=25,\theta_{1}=\frac{\pi}{12},t_{1}=4),(s_{2}=25,\theta_{2}=\frac{\pi}{12},t_{2}=2),(s_{3}=25,\theta_{3}=\frac{\pi}{18})$.